# Neural Ordinary Differential Equations for Learning and Extrapolating System Dynamics Across Bifurcations

### v1

@@ Abstract
Forecasting system behaviour near and across bifurcations is crucial for identifying potential shifts in dynamical systems. While machine learning has recently been used to learn critical transitions and bifurcation structures from data, most studies remain limited as they exclusively focus on discrete-time methods and local bifurcations. To address these limitations, we use Neural Ordinary Differential Equations which provide a continuous, data-driven framework for learning system dynamics. We apply our approach to a predator-prey system that features both local and global bifurcations, presenting a challenging test case. Our results show that Neural Ordinary Differential Equations can recover underlying bifurcation structures directly from timeseries data by learning parameter-dependent vector fields. Notably, we demonstrate that Neural Ordinary Differential Equations can forecast bifurcations even beyond the parameter regions represented in the training data. We also assess the method’s performance under limited and noisy data conditions, finding that model accuracy depends more on the quality of information that can be inferred from the training data, than on the amount of data available.
@@ 1 Introduction
Bifurcation theory is a powerful tool to study how and when a system undergoes sudden shifts in its dynamical behaviour. A bifurcation is a point where a small change in the system’s parameters or external drivers cause a qualitative shift in its dynamical behaviour. Because these shifts can have significant consequences, ranging from smaller-scale ecosystem collapses [1, 2, 3, 4] to large-scale climate tipping points [5, 6, 7], understanding and predicting system behaviour near bifurcation points is crucial [8]. Identifying the types of bifurcations that a real-world system can exhibit provides insights into the system’s long-term dynamics and can help guide interventions to prevent undesired outcomes.
@@
A major challenge lies in detecting bifurcations in systems where the underlying processes and governing equations are poorly understood or unknown. In response, data-driven approaches have gained increasing interest. In particular, machine learning is being actively explored as a tool for capturing system dynamics and bifurcations directly from data [9, 10]. Deep learning methods, for instance, outperform classical early warning indicators in both sensitivity and specificity when detecting pre-bifurcation signals in time series data[11, 12, 13]. Parameter-driven reservoir computing has successfully extrapolated the dynamics of bifurcating systems and, in some cases, even predicted behaviour beyond the training data [14, 15, 16, 17, 18]. More recently, unsupervised learning techniques, which do not require prior knowledge of the bifurcation parameter, have been shown to infer the parameter’s value directly from data [19]. Additionally, hybrid modelling, which combines mechanistic differential equations with data-driven approximators, such as neural networks, has proven to effectively capture the bifurcation behaviour of diverse systems[20, 21].
@@
Despite these advances, most existing machine learning applications to bifurcations remain limited. Many focus exclusively on discrete machine learning methods [11, 12, 13, 19, 14, 15, 16, 18], which rely on discrete-time jumps to capture dynamics. In contrast, many real-world systems evolve smoothly with gradual shifts rather than discrete steps. This mismatch between the discrete assumptions of most machine learning methods and the continuous nature of many systems limits their effectiveness in detecting bifurcations in continuous-time settings. Moreover, much of the focus has been on local bifurcations [11, 12, 20, 21], which can be identified from local properties of the system. In contrast, global bifurcations depend on the structure of the large-scale state space. Because global bifurcations can have significant consequences for system dynamics[22], it is important to investigate methods that can infer global system structure from limited data and thereby detect this class of bifurcations.
@@
To address these limitations, we use Neural Ordinary Differential Equations (ODEs), a class of models introduced by Chen et al.[23] that use neural networks to represent continuous-time dynamics, serving as a data-driven alternative to traditional ODEs. Instead of relying on explicitly defined equations, Neural ODEs can be used to learn a continuous vector field directly from data. This makes them well-suited for modelling complex, non-linear, or unknown systems[23, 24, 25, 26]. The continuous vector fields learned by Neural ODEs offer a promising approach for uncovering bifurcation structures from data, as they can capture smooth changes in the state space that may lead to qualitative shifts in system behaviour [24, 27].
@@
In this work, we demonstrate the potential of Neural ODEs as a tool for extrapolating bifurcation dynamics from data. Although Neural ODEs, and even parameterized modifications[28], are established techniques, our approach introduces a novel application: explicitly incorporating a bifurcation parameter as an input to learn bifurcation structures. This specific adaptation enables the model to capture and analyse regime shifts, offering valuable insights into system behaviour. By directly learning the underlying vector field, Neural ODEs provide a global continuous-time representation of the system dynamics, making them particularly well-suited for capturing not only local bifurcations, but also global ones. We study a predator-prey system, that exhibits both local and global bifurcations, making it a challenging test case. We investigate the ability of Neural ODEs to extrapolate dynamics across bifurcations. To examine their robustness, we conduct experiments with limited trajectory data and added measurement noise, demonstrating their potential as a tool for predicting system transitions.
@@
The rest of the paper is organized as follows. In Sec.  2, we give a brief overview on Neural ODEs and how they can be adapted to include bifurcation parameters. In Sec. 3, we introduce our predator-prey test case and summarise its bifurcation structure. In Sec. 4, we discuss the set-up of our simulation experiments. In Sec.  5, we showcase our findings, which will be further discussed in Sec. 6.
@@ 2 Neural ODEs with bifurcation parameter
Neural ODEs, introduced by Chen et al.[23], provide a framework for learning and describing continuous time dynamics using neural networks. Like classical ODEs they model system evolution through a differential equation:
@@
Throughout this paper, we use the term model specifically to refer to the Neural ODE architecture, and model parameters to refer to its internal weights and biases. We will use the term system when referring to a dynamical system and its corresponding equations.
@@
where $\alpha\in\mathbb{R}^{m}$ are the bifurcation parameters or external drivers. Note that in our formulation we omit $t$ as input, since we assume for simplicity that our system is autonomous.
@@
Representing system dynamics as a vector field allows us to simulate how the system evolves at any point in time, from any initial condition. In contrast, commonly used deep learning methods for time series forecasting, such as Recurrent Neural Networks, learn discrete-time mappings from one time point to the next, which limits them to fixed time intervals[24, 29, 30]. A continuous representation of dynamics, as used in Neural ODEs, allows us to analyse qualitative system behaviour, such as bifurcations, using approaches similar to those used in the analysis of classical ODEs.
@@
Neural ODEs compute trajectories by iteratively integrating the learned dynamics forward in time (Fig. 1). At each forward step, the network receives the current state $z_{t^{*}}$ and bifurcation parameter $\alpha$ as inputs and outputs the corresponding time derivative of the state $\frac{dz}{dt}$. The outputted derivative together with the current state is integrated using an ODE solver, which then produces the state variables at the next time point $z_{t^{*}+\delta t}$. This iterative process allows Neural ODEs to generate trajectories that capture the underlying system dynamics. Note that we are not limited to fixed time steps, since the ODE solver can take arbitrarily small or adaptive steps to closely approximate the continuous dynamics of the system.
During training, we optimize the model parameters $\theta$ based on time series data from our system for different values of $\alpha$. During each step in the optimisation process, trajectories are generated from initial conditions, as described above. The loss is computed by comparing the predicted and true states at the observed data points. The model parameters are then updated using backpropagation, which requires computing the gradient of the loss with respect to each of the model parameters.
3 Test case: predator-prey system with local and global bifurcation
To test the bifurcation parameter-dependent Neural ODE, we applied it to a non-linear predator-prey system. In particular, the interaction between prey and predator is modelled by the following:
$\begin{split}\frac{dx}{dt}&=3x(1-x)-xy-\alpha(1-e^{-5x}),\\ \frac{dy}{dt}&=-y+3xy,\end{split}$ (3)
where $x$ and $y$ represent the prey and predator populations, respectively, and $\alpha\in\mathbb{R}^{1}$ serves as an external driver and bifurcation parameter. This system of equations represents a modification of the classic Lotka–Volterra equations with logistic growth for the prey population. The final term in Eq. 3 represents a non-linear damping effect on the prey population. This term increases with both the prey population size $x$ and the bifurcation parameter $\alpha$, indicating that higher values of $x$ and $\alpha$ lead to a larger reduction in the prey’s growth rate.
This system presents an ideal test case since the non-linear and non-polynomial dampen term allow us to evaluate the ability of the neural ODE to handle more complex dynamics. Additionally, in the parameter region we study, the system exhibits two distinct types of bifurcations: a local Hopf bifurcation and a global homoclinic bifurcation. Successfully capturing both types of bifurcations requires the model to infer both local stability changes and global transitions in state-space behaviour.
[…]
@@ 3 Test case: predator-prey system with local and global bifurcation
Figure 3: From trajectory data to bifurcation diagram. (a) Trajectory data is used as input to train the neural ODE. (b) The Neural ODE defines a learned vector field, which can be directly extracted from the model. (c) We determine the bifurcation structure from the vector field (Appendix C).
@@
Fig. 2 illustrates the bifurcation structure of the system, highlighting the different dynamical regimes. Note that we limit ourself to the bifurcation parameter range $\alpha\in(0.6,0.8)$, since the zero steady state is only stable for $\alpha>0.6$. Within this range the system exhibits three distinct dynamical regimes, which allows us to analyse if the model is able to extrapolate transitions between qualitatively different behaviours. The system has four steady states: the zero steady state, two saddle nodes on the $x$-axis and a coexistence steady state. In Regime A the positive steady state is stable, meaning that for certain initial conditions predator and prey populations settle into a steady coexistence. At $\alpha=0.67$, the system undergoes a Hopf bifurcation, causing the coexistence steady state to lose stability, giving rise to stable limit cycles, which characterise Regime B. These limit cycles persist until $\alpha\approx 0.71$, where the system undergoes a global heteroclinic bifurcation; the limit cycle collides with two saddle nodes, leading to a collapse. In Regime C ($\alpha>0.71)$, the only remaining stable equilibrium is the zero steady state, meaning that both populations eventually go extinct, marking a critical transition point, that has previously been interpreted as an ecological phenomenon referred to as over-exploitation[31].
Extensions of Neural ODEs have previously been applied to more basic forms of predator-prey systems, such as the classical Lotka-Volterra system [32, 33, 34]. However, these works focused on learning dynamics from time series data with fixed parameters, using trajectories within a single dynamical regime. Consequently, they did not address how system behaviour changes with a bifurcation parameter or explore predictions across qualitatively different dynamical regimes.
4 Experiment set-up
We designed a series of simulation experiments to evaluate how well the Neural ODE captures system dynamics and generalises across bifurcations. Using simulated trajectories from the predator-prey system, we investigated whether the model can reconstruct the underlying vector field and bifurcation structure from time series data (Fig. 3).
@@ 4 Experiment set-up
A primary dataset was used for two purposes: first, to tune the model’s hyperparameters (see Appendix B for more details), and second, to showcase that the Neural ODE can infer the full vector field and bifurcation structure from limited trajectory data. In addition to this primary experiment, we conducted targeted experiments to explore how specific factors, such as data size and measurement noise, affect the model’s ability to learn the underlying dynamics. Below, we summarize the details of the different experiments (Fig. 4). For a more detailed overview see Appendix A.
@@
  • Primary experiment. The data for the primary experiment consisted of trajectories sampled at five different values of the bifurcation parameter $\alpha$, with two distinct initial conditions for each value, yielding a total of ten time series. Sampling was restricted to regime A and B (as previously defined in Fig. 2), enabling us to assess whether the Neural ODE could extrapolate the global bifurcation and collapse in regime C.
@@
  • Experiment 1: Extrapolation across different regimes. We trained the Neural ODE on trajectories from specific dynamical regimes of the bifurcation diagram to assess its ability to extrapolate to unseen regimes. In one case, the training data was drawn from both regime A and B (the primary dataset). In another, training was restricted to the smaller regime B, which exhibits only limit cycles. Lastly, we trained the model exclusively on data from regime A, where only a stable focus is present.
@@
  • Experiment 2: Limited data. We investigated how the amount of training data influences the model performance. To explore this, we created two datasets, for which we reduced the available data compared to the primary dataset, either by decreasing the number of sampled $\alpha$-values or by computing trajectories for only a single initial condition.
@@
    where $\hat{x}_{i}$, $\hat{y}_{i}$ are our noisy data points and $\sigma$ controls the noise level. We considered three cases: no noise $(\sigma=0)$, low noise $(\sigma=0.05)$, and high noise $(\sigma=0.2)$. Using these values of $\sigma$, the noisy data we generated remained positive for both $x$ and $y$.
Figure 4: Sampling locations of training data for different experiments indicated by dashed lines. Experiment 1: Data is sampled from different regimes. Experiment 2: Amount of data is limited, either by decreasing the number of sampled $\alpha$-values or by reducing the number of initial conditions from two to one (dotted lines). Experiment 3: Different levels of measurement noise are added to the data.
The neural network architecture, ODE solver settings and hyperparameters remained fixed across all experiments. Hyperparameters were selected based on tuning performed on the primary dataset (see Appendix B for more details). For our architecture we used a Multilayer Perceptron (MLP) coupled with an adaptive ODE solver (DOPRI5[35]) for integration. The number of layers and depth of each layer were varied during the tuning process to explore different model complexities. We use the torchdiffeq library [36], which provides ODE solvers implemented in PyTorch. Backpropagation can be performed either by differentiating through all intermediate solver steps or by using the adjoint sensitivity method[23], which reduces memory usage during training. We used standard backpropagation through the internals of the solver, and did not encounter memory or stability issues.
To guide the training process, we combined a mean absolute error (MAE) with a physics-informed term to obtain our loss function:
$\mathcal{L}=\frac{1}{N}\sum_{i=1}^{N}{\|\mathbf{z}_{i}-\tilde{\mathbf{z}}_{i}\|}_{1}+\lambda\left[\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{x}}{dt}(0,y_{j})\right|+\right.\left.\frac{1}{M}\sum_{j=1}^{M}\left|\frac{d\tilde{y}}{dt}(x_{j},0)\right|\right],$
where $\tilde{\mathbf{z}}_{i}=[\tilde{x}_{i},\tilde{y}_{i}]$ are our predicted state variables, $N$ is the number of data points and $M$ indicates the number of points on which we evaluate the physics-loss. Unlike mean squared error (MSE), which disproportionately emphasizes large errors, MAE treats all deviations linearly[37, 38]. Since MSE would assign greater weight to outliers and might obscure small but meaningful errors during training, MAE is a more suitable objective function for our application.
The physics part of the loss function enforces that the axes are invariant by penalizing non-zero derivatives of the system along specific axes. Since Neural ODEs model derivatives directly, and much of our prior knowledge about physical systems is expressed through differential equations, it is natural to incorporate physics-based information into the loss function. Such physics-informed neural ODEs have been explored in previous work [39, 40, 41]. In our experiments, adding the physics-informed term improved generalisation to unseen parameter ranges by guiding the model training in regions of the vector field not covered by the data. In Equation 4, the hyperparameter $\lambda$ is the physics-loss weight and determines how much the physics-loss contributes to the total loss. Note that the physics-informed term is not a hard constraint, but a soft constraint driving the derivatives on the axes towards zero.

### v2

@@ Abstract
Forecasting system behaviour near and across bifurcations is crucial for identifying potential shifts in dynamical systems. While machine learning has recently been used to learn critical transitions and bifurcation structures from data, most studies remain limited as they exclusively focus on discrete-time methods and local bifurcations. To address these limitations, we use Neural Ordinary Differential Equations which provide a data-driven framework for learning system dynamics. Our results show that Neural Ordinary Differential Equations can recover underlying bifurcation structures directly from time-series data by learning parameter-dependent vector fields. Notably, we demonstrate that Neural Ordinary Differential Equations can forecast bifurcations even beyond the parameter regions represented in the training data. We demonstrate our approach on three test cases: the Lorenz system transitioning from non-chaotic to chaotic behaviour, the Rössler system moving from chaos to period doubling, and a predator-prey model exhibiting collapse via a global bifurcation.
@@ 1 Introduction
Bifurcation theory is a powerful tool to study how and when a system undergoes sudden shifts in its dynamical behaviour. A bifurcation is a point where a small change in the system’s parameters or external drivers cause a qualitative shift in its dynamical behaviour. Because such shifts can have significant consequences, ranging from smaller-scale ecosystem collapses [1, 2, 3, 4] to large-scale climate tipping points [5, 6, 7], understanding and predicting system behaviour near bifurcation points is crucial [8]. Identifying the types of bifurcations that a real-world system can exhibit provides insights into the system’s long-term dynamics and can help guide interventions to prevent undesired outcomes.
@@
A major challenge lies in detecting bifurcations in systems where the underlying processes and governing equations are poorly understood or unknown. In response, data-driven approaches have gained increasing interest. In particular, machine learning is being actively explored as a tool for capturing system dynamics and bifurcations directly from data [9, 10]. Deep learning methods, for instance, outperform classical early warning indicators in both sensitivity and specificity when detecting pre-bifurcation signals in time series data[11, 12, 13]. Parameter-driven reservoir computing has successfully extrapolated the dynamics of bifurcating systems and, in some cases, even predicted behaviour beyond the training data [14, 15, 16, 17, 18]. More recently, unsupervised learning techniques, which do not require prior knowledge of the bifurcation parameters, have been shown to infer the parameter values directly from data [19]. Additionally, hybrid modelling, which combines mechanistic modelling with data-driven approximators, such as neural networks, has proven to effectively capture the bifurcation behaviour of diverse systems[20, 21].
@@
Despite these advances, most existing machine learning applications to bifurcations remain limited. Many focus exclusively on discrete machine learning methods [11, 12, 13, 19, 14, 15, 16, 18], which rely on discrete-time jumps to capture dynamics. In contrast, many real-world systems evolve smoothly with gradual shifts rather than discrete steps. This mismatch between the discrete assumptions of most machine learning methods and the continuous nature of many systems limits their effectiveness in detecting bifurcations in continuous-time settings. Moreover, much of the focus has been on local bifurcations [11, 12, 20, 21], which can be identified from local properties of the system. In contrast, global bifurcations depend on the structure of the large-scale state-space. Because global bifurcations can have significant consequences for system dynamics[22], it is important to investigate methods that can infer global system structure from limited data and thereby detect this class of bifurcations.
@@
To address these limitations, we use Neural Ordinary Differential Equations (ODEs); a class of models introduced by Chen et al.[23] that use neural networks to approximate continuous-time dynamics, serving as a data-driven alternative to traditional ODEs. Instead of relying on explicitly defined equations, Neural ODEs can be used to learn a vector field directly from data. This makes them well-suited for modelling complex, nonlinear, or unknown systems[23, 24, 25, 26]. The vector fields learned by Neural ODEs offer a promising approach for uncovering bifurcation structures from data, as they can capture smooth changes in the state-space that may lead to qualitative shifts in system behaviour [24, 27].
@@
In this work, we demonstrate the potential of Neural ODEs as a tool for extrapolating bifurcation dynamics from data. Although Neural ODEs, and even parameterized modifications[28], are established techniques, our approach introduces a novel application: explicitly incorporating a bifurcation parameter as an input to learn bifurcation structures. This specific adaptation enables the model to capture and analyse regime shifts, offering valuable insights into system behaviour. We examine three test cases: the Lorenz system transitioning to chaotic behaviour, the Rössler system moving from chaos to period-doubling and a predator–prey system exhibiting both local and global bifurcations. For each of these systems we investigate the ability of Neural ODEs to extrapolate dynamics across bifurcations. To examine their robustness, we conduct additional experiments on the predator-prey model with limited trajectory data and added measurement noise.
@@
The rest of the paper is organized as follows. In Sec.  2, we give a brief overview on Neural ODEs and how they can be adapted to include bifurcation parameters. In Sec. 3, we introduce our different test cases and discuss the training procedure. In Sec.  4, we showcase our findings, which will be further discussed in Sec. 5.
@@ 2 Neural ODEs with bifurcation parameter
Neural ODEs provide a framework for learning and describing continuous-time dynamics using neural networks[23]. Like classical ODEs they model system evolution through a differential equation:
@@
Throughout this paper, we use the term model specifically to refer to the Neural ODE architecture, and model parameters to refer to its internal weights and biases. We will use the terms system and system parameters when referring to a dynamical system and its corresponding equations.
@@
where $\alpha\in\mathbb{R}^{m}$ are the bifurcation parameters. Note that in our formulation we omit $t$ as input, since we assume for simplicity that our system is autonomous.
@@
Figure 2: From trajectory data to bifurcation diagram. (a) Trajectory data is used as input to train the Neural ODE. (b) The Neural ODE defines a learned vector field, which can be directly extracted from the model. (c) We determine the bifurcation structure from the vector field (Appendix D).
Representing system dynamics as a vector field allows us to simulate how the system evolves at any point in time, from any initial condition. While any computer implementation must approximate solutions using floating-point arithmetic and discrete solver steps, a Neural ODE defines a numerical approximation to continuous-time vector fields rather than an inherently discrete-time mapping. In contrast, commonly used deep learning methods for time series forecasting, such as Recurrent Neural Networks, learn discrete mappings from one time point to the next, which limits them to fixed time intervals[24, 29, 30]. A continuous representation of dynamics, as used in Neural ODEs, allows us to analyse qualitative system behaviour, such as bifurcations, using approaches similar to those used in the analysis of classical ODEs.
Neural ODEs compute trajectories by iteratively integrating the learned dynamics forward in time (Fig. 1). At each forward step, the network receives the current state $z_{t^{*}}$ and bifurcation parameter $\alpha$ as inputs and outputs the corresponding time derivative of the state $dz/dt$. The outputted derivative together with the current state is integrated using an ODE solver, which then produces the state variables at the next time point $z_{t^{*}+\delta t}$. This iterative process allows Neural ODEs to generate trajectories that capture the underlying system dynamics. Note that we are not limited to fixed time steps, since the ODE solver can take arbitrarily small or adaptive steps to closely approximate the continuous dynamics of the system.
During training, we optimize the model parameters $\theta$ based on time series data from our system for different values of the bifurcation parameter $\alpha$. During each step in the optimisation process, trajectories are generated from initial conditions, as described above. The loss is computed by comparing the predicted and true states at the observed data points. The model parameters are then updated using backpropagation, which requires computing the gradient of the loss with respect to each of the model parameters.
In this study, we examine multiple systems to determine whether a Neural ODE can learn bifurcation structures directly from data. We follow the same approach in each case: train a Neural ODE model on time-series data and reconstruct the bifurcation structure from the learned dynamics (Fig. 2). More specifically, we train the model on a particular dynamical regime and investigate whether it can extrapolate to different dynamical behaviours across bifurcations.
3 Experimental setup
To test the bifurcation parameter–dependent Neural ODE, we applied it to three test cases. This section describes the different systems and explains the experimental setup that was used. For the first two test cases, the Lorenz and Rössler systems, we conducted a single experiment to showcase the ability of the Neural ODE to infer the bifurcation structure from limited trajectory data. For the third test case, a predator–prey model, we perform additional experiments to investigate how factors such as data size and measurement noise influence the model’s ability to capture the underlying dynamics. Fig. 3 shows the bifurcation structures of the different systems and highlights the parameter ranges used for training and extrapolation. Additional details on the generation of the synthetic data can be found in Appendix B.
Figure 3: Bifurcation diagrams and sampling locations of training data for different systems and experiments. Top: Lorenz system. Middle: Rössler system. Bottom: Predator-prey system; Experiment 1: Data is sampled from different regimes; Experiment 2: Amount of data is limited, either by decreasing the number of sampled $\alpha$-values or by reducing the number of initial conditions from two to one; Experiment 3: Different levels of measurement noise are added to the data. Line style indicates amount of initial conditions (striped = two initial conditions, dotted = one initial condition).
Test case I: Chaos in the Lorenz system
For our first test case, we considered the classic Lorenz system:
$\begin{split}\dot{x}&=\sigma(y-x),\\ \dot{y}&=rx-y-xz,\\ \dot{z}&=xy-bz,\end{split}$ (3)
with fixed parameters $\sigma=10$ and $b=8/3$, and $r$ serving as the bifurcation parameter. Bifurcations and chaos in the Lorenz system have extensively been studied[31]. We focused on the transition from non-chaotic to chaotic behaviour as $r$ is varied. More specifically, we investigated the parameter range $r\in(1,30)$}, where the Lorenz system becomes chaotic near $r\approx 24.74$. The data consisted of trajectories sampled from the non-chaotic regime at eight different values of the bifurcation parameter $r$ with two distinct initial conditions for each value, yielding a total of sixteen time series.
Test case II: From chaos to period-doubling in the Rössler system
For our second test case, we considered the Rössler system:
$\begin{aligned} \dot{x}&=-y-z,\\ \dot{y}&=x+ay,\\ \dot{z}&=b+z(x-c),\end{aligned}$ (4)
with parameters $a=b=1/5$ fixed, while treating $c$ as the varying bifurcation parameter. The Rössler system is another well-studied chaotic system [32, 33, 34]. Contrary to the previous test case, we trained our model in the chaotic regime and attempted to extrapolate to lower values of $c$ where the system exhibits limit cycles undergoing period-doubling bifurcations. More specifically, we focused on the parameter range $c\in(3,6)$}. The training data for this test case consisted of ten trajectories sampled at ten different values of the bifurcation parameter $c\in(4.5,5.5)$ with only one initial condition for each value.
Test case III: Global bifurcation in predator-prey system
For our third test case, we considered the interaction between prey and predator, modelled by the following:
$\begin{split}\dot{x}&=3x(1-x)-xy-\alpha(1-e^{-5x}),\\ \dot{y}&=-y+3xy,\end{split}$ (5)
where $x$ and $y$ represent the prey and predator populations, respectively, and $\alpha$ serves as the bifurcation parameter. This system of equations extends the classic Lotka–Volterra equations with logistic growth for the prey and an additional nonlinear harvesting term of Ivlev type. This term makes that higher values of the bifurcation parameter $\alpha$, reduce prey growth at higher densities. While this exact formulation has not been studied before, closely related predator–prey models with Ivlev-type responses have been analyzed in the literature[35, 36, 37, 38], and we provide a detailed analysis of the present system in Appendix A.
@@ 3 Experimental setup
We limit ourself to the bifurcation parameter range $\alpha\in(0.6,0.8)$, where the system exhibits two distinct types of bifurcations: a local Hopf bifurcation ($\alpha=0.67$) and a global heteroclinic bifurcation ($\alpha\approx 0.71$). We consider three different dynamical regimes. In Regime A $(0.6,0.67)$ there exists a stable positive steady state, meaning that for a range of initial conditions the predator and prey populations settle into a steady coexistence. Regime B $(0.67,0.71)$ is characterised by limit cycles that persist until Regime C ($\alpha>0.71)$, where the only remaining stable equilibrium is the zero steady state, meaning that both populations eventually go extinct.
@@
For this test case, we designed a series of simulation experiments to evaluate how well the Neural ODE captures system dynamics and generalises across bifurcations. Similar to the other test cases, a primary dataset was used to showcase that the Neural ODE can infer the full vector field and bifurcation structure from limited trajectory data. In addition, we conducted more targeted experiments to explore how specific factors, such as data size and measurement noise, affect the model’s ability to learn the underlying dynamics. Below, we summarize the details of the different experiments.
@@
  • Primary experiment. The data for the primary experiment consisted of trajectories sampled at five different values of the bifurcation parameter $\alpha$, with two distinct initial conditions for each value, yielding a total of ten time series. Sampling was restricted to regime A and B, enabling us to assess whether the Neural ODE could extrapolate the global bifurcation and collapse in regime C.
@@
  • Experiment 1: Extrapolation across different regimes. We trained the Neural ODE on trajectories from specific dynamical regimes of the bifurcation diagram to assess its ability to extrapolate to unseen regimes. In one case, the training data was drawn from both regime A and B (the primary dataset). In another, training was restricted to the smaller regime B, which exhibits only limit cycles. Lastly, we trained the model exclusively on data from regime A, where the positive steady state is stable.
@@
  • Experiment 2: Limited data. We investigated how the amount of training data influences the model performance. To explore this, we created two datasets for which we reduced the available data compared to the primary dataset, either by decreasing the number of sampled $\alpha$-values or by computing trajectories for only a single initial condition.
@@
    where $\hat{x}_{i}$, $\hat{y}_{i}$ are our noisy data points and $\sigma$ controls the noise level. We considered three cases: no noise $(\sigma=0$; i.e., the primary dataset), low noise $(\sigma=0.05)$, and high noise $(\sigma=0.2)$. Using these values of $\sigma$, the noisy data we generated remained positive for both $x$ and $y$.
Neural ODE Architecture and Training Procedure
For each of the systems we used a similar architecture: a Multilayer Perceptron (MLP) coupled with an ODE solver for time integration. For the Lorenz and Rössler systems we used a fixed-step fourth order Runge-Kutta solver. For the predator–prey system we instead employed the adaptive solver DOPRI5 in order to achieve higher precision. We used the torchdiffeq library [39], which provides ODE solvers implemented in PyTorch. Hyperparameters were selected based on tuning performed on a validation set (See Apppendix C.3 for more details). The number of layers and depth of each layer were varied during the tuning process to explore different model complexities. Backpropagation can be performed either by differentiating through all intermediate solver steps or by using the adjoint sensitivity method[23], which reduces memory usage during training. We used standard backpropagation through the internals of the solver, and did not encounter memory or stability issues.
To guide the training process, we used a mean absolute error (MAE) for the loss function. Compared to the mean squared error (MSE), which disproportionately emphasizes large errors, MAE treats all deviations linearly[40, 41]. Since MSE would assign greater weight to outliers and might obscure small but meaningful errors during training, MAE is a more suitable objective function for our application.
It is important to note that for the predator-prey model we additionally added a physics part to the loss function, which enforces that the axes are invariant by penalizing non-zero derivatives of the system along specific axes (See Appendix C.1). Since Neural ODEs model derivatives directly, and much of our prior knowledge about physical systems is expressed through differential equations, it is natural to incorporate physics-based information into the loss function. Such physics-informed Neural ODEs have been explored in previous work [42, 43, 44]. In our experiments, adding the physics-informed term improved generalisation to unseen parameter ranges by guiding the model training in regions of the vector field not covered by the data.